# Spatial super-resolution in nanosensing with blinking emitters

### v1

@@ IV Discussion
We believe that the proposed approach and its proof-of-concept experiment can pave the way for:
@@
  • extension of the approach to multimodal sensing, when several characterisics of the sample are reconstructed simultaneously;
@@
V Conclusions
We proposed a simple and elegant technique capable of combining sensing with super-resolution imaging, based on non-trivial temporal statistics typical for blinking fluorescent nanoemitters. The approach is universal since the dependence of the fluorophores response on the characteristic to be reconstructed can take different form: changes of the spectrum, fluorescence lifetime, etc. It is shown that splitting the fluorescence signal into just two detection channels at each spatial detection point is sufficient for reliable reconstruction of the local characteristic distribution.
The conducted model experiment proofs soundness and efficiency of the approach and can be used as a guidance for the real-world experiments with, for example, living cells and structured fluorescent nanodiamonds [48] or blinking graphene quantum dots ensembles [49]. In agreement with the theoretical predictions, usage of cumulants instead of emission intensity enhances sharpness of the reconstructed parameter distribution and enables analysis of smaller features of the sample.
Proposed methodology will be useful for local measurements of living cells’ physical parameters, such as temperature, pH, electric field gradient, presence of reactive oxygen spices, i.e. the characteristics which could be recognized as spectral changes of appropriate fluorophore delivered targetly to the wanted cell or even organelle. In its turn, this advanced local sensing tool could give an important insights into fundamental life science, which still lacks instruments capable of monitoring the cellular balance with nanometre spatial resolution to guarantee measurements consistency for that dynamic system.
@@ Acknowledgements.
This study was accomplished through the financial support of Research Council of Finland (Flagship Programme PREIN, decision 320166; research project, decision 357033); QuantERA project EXTRASENS, in Finland funded through Research Council of Finland, decision 361115), Horizon Europe MSCA FLORIN Project 101086142.
Appendix A Optimal estimator for linear-response sensing
Here, we show that the information about the environment parameter $\theta$, contained in the full signal density profile $\varepsilon(x,\theta)$, described by Eq. (11), can be efficiently compressed into two scalar characteristics $Y_{i}$ (Eq. (14)), which are used for building the estimator $\bar{\theta}$ according to Eq. (12). The analysis is performed in the vicinity of the reference value $\theta_{0}$ of the parameter, i.e. the deviation $|\theta-\theta_{0}|$ is assumed to be small.
We assume that the signal profile is measured at a discrete set of bins $\{x\}$ and contains shot noise satisfying Poisson statistics:
$\varepsilon(x,\theta)=S_{0}(x)+S_{1}(x)(\theta-\theta_{0})+\delta(x),$ (36)
where $\operatorname{E}[\delta(x)]=0$, $\operatorname{E}[\delta(x)\delta(x^{\prime})]\approx S_{0}(x)\delta_{xx^{\prime}}$. To reduce the influence of the noise caused by the emitters’ blinking on the final estimates of the environment parameter, it is useful to consider the normalized signal $\varepsilon(x,\theta)/\sum_{x^{\prime}}\varepsilon(x^{\prime},\theta)$. The Fisher information of that normalized density profile relatively to the parameter $\theta$ at the point $\theta=\theta_{0}$ can be calculated as
$F_{\text{full}}(\theta)=\sum_{x}\frac{S_{1}^{2}(x)}{S_{0}(x)}-\frac{1}{\sum_{x}S_{0}(x)}\Bigl{(}\sum_{x}S_{1}(x)\Bigr{)}^{2},$ (37)
where the second term describes “information loss” caused by the normalization [50].
For a low level of noise, the estimator $\bar{\theta}$ (constructed from the noisy signal according to Eq. (12)) is a random variable obeying approximately normal distribution with $\operatorname{E}[\bar{\theta}]=\theta$ and $\operatorname{Var}[\bar{\theta}]\approx\sum_{x}Q_{1}^{2}(x)S_{0}(x)$. The Fisher information after “compression” of the profile $\varepsilon(x,\theta)$ into the scalar $\bar{\theta}$ equals
$F_{\bar{\theta}}(\theta)=\Bigl{[}\sum_{x}Q_{1}^{2}(x)S_{0}(x)\Bigr{]}^{-1}.$ (38)
According to general properties of Fisher information [51, 52], the following inequality holds:
$F_{\bar{\theta}}(\theta)\leq F_{\text{full}}(\theta).$ (39)
On the other hand, one can prove by direct calculation that the choice
$Q_{i}(x)=\frac{q_{i}(x)}{\sum_{x^{\prime}}q_{i}(x^{\prime})S_{i}(x^{\prime})},\quad i=0,1,$ (40)
@@ Appendix A Optimal estimator for linear-response sensing
$q_{0}(x)=S_{0}(x)-S_{1}(x)\frac{\sum_{x^{\prime}}S_{0}(x^{\prime})S_{1}(x^{\prime})}{\sum_{x^{\prime}}S_{1}^{2}(x^{\prime})}$ (41)
@@
$q_{1}(x)=\frac{S_{1}(x)}{S_{0}(x)}-\frac{\sum_{x^{\prime}}S_{1}(x^{\prime})}{\sum_{x^{\prime}}S_{0}(x^{\prime})}$ (42)
ensures that the maximal value of Fisher information is reached: $F_{\bar{\theta}}(\theta)=F_{\text{full}}(\theta)$.
Appendix B Practical estimator for sensing with two spectral or temporal windows
In this appendix, we show how scalar characteristics $Y_{i}$, defined by Eq. (14) without any particular constraints imposed on the weights $Q_{j}(x)$, can be modified so that Eq. (13) are satisfied.
In general case, the values $Y_{i}$ are decomposed according to Eq. (17). One can construct the following linear combinations of those quantities:
$\bar{Y}_{0}=\frac{Y_{0}A_{11}-Y_{1}A_{01}}{A_{00}A_{11}-A_{10}A_{01}}\mapsto 1,$ (43)
$\bar{Y}_{1}=\frac{Y_{1}A_{00}-Y_{0}A_{10}}{A_{00}A_{11}-A_{10}A_{01}}\mapsto\theta-\theta_{0}.$ (44)
Linearity of Eq. (17) implies that the weights $\bar{Q}_{i}(x)$, corresponding to the quantities $\bar{Y}_{j}$ are related to the initial weights $Q_{i}(x)$ by the same relations:
$\bar{Q}_{0}(x)=\frac{Q_{0}(x)A_{11}-Q_{1}(x)A_{01}}{A_{00}A_{11}-A_{10}A_{01}},$ (45)
$\bar{Q}_{1}(x)=\frac{Q_{1}(x)A_{00}-Q_{0}(x)A_{10}}{A_{00}A_{11}-A_{10}A_{01}}.$ (46)
One can check by direct substitution that the conditions (13) are satisfied for the constructed weights $\bar{Q}_{i}(x)$.
Fig. 9 shows an example of weights calculated for the model signal densities depicted in Fig. 1(a,b).
Figure 9: Weights for sensing by fluorescence spectrum (a,c,e) and lifetime (b,d,f) for the model signal densities depicted in Fig. 1(a,b): optimal weights (a,b) defined by Eq. (40); practical weights (c,d) corresponding to Gaussian spectral filters and threshold-based treatment of fluorescence lifetime; optimal linear combinations (e,f) of practical weights accoring to Eqs. (45) and (46).
@@ Appendix B Practical estimator for sensing with two spectral or temporal windows
Appendix C General properties of cumulant-based estimator for noiseless data
@@
To be physically reasonable, the estimator $\bar{\theta}_{\text{nonlin}}(\mathbf{r})$, defined by Eq. (26) should satisfy several general requirements when applied to noiseless data.
1. For spatially uniform distribution $\theta(\mathbf{r})=const=\theta$ the estimator should reconstruct the correct value: $\bar{\theta}_{\text{nonlin}}(\mathbf{r})=\theta$ for all $\mathbf{r}$. In that case, according to Eq. (22), one has $C_{1}^{(n)}(\mathbf{r})/C_{0}^{(n)}(\mathbf{r})=Y_{1}^{n}/Y_{0}^{n}$ and, therefore, Eq. (24) retains its value after the replacement $C_{j}^{(n)}(\mathbf{r})\mapsto Y_{j}^{n}$. Then, according to the definition (26), the equality $\bar{\theta}_{\text{nonlin}}(\mathbf{r})=\theta$ holds for all $\mathbf{r}$.
2. If the parameter values $\theta_{k}$ in the vicinity of the emitters are bounded as $\theta_{\text{min}}\leq\theta_{k}\leq\theta_{\text{max}}$, the estimator value is also bounded: $\theta_{\text{min}}\leq\bar{\theta}_{\text{nonlin}}(\mathbf{r})\leq\theta_{\text{max}}$ for all $\mathbf{r}$. According to Eq. (15), the estimator $\bar{\theta}=\Theta(\theta)$ is a function of the ratio $Y_{1}/Y_{0}$, which itself is a function of $\theta$. The monotonicity of the dependence $\Theta(\theta)$ implies that both of the above-mentioned functions are also monotonic. Therefore, for each emitter, the ratio $Y_{k1}/Y_{k0}$ is bounded by its values at $\theta_{k}=\theta_{\text{min}}$ and $\theta_{k}=\theta_{\text{max}}$. The same conclusion is valid for the ratio $Y_{k1}^{n}/Y_{k0}^{n}$, since $n$ is positive.
Assuming that the blinking statistics is the same for all emitters (i.e., $C_{t}^{(n)}[s_{k}(t)]$ depends only on $n$, but not on $k$) and taking into account the definition (22), one can conclude that the ratio $R(\mathbf{r})=C_{1}^{(n)}(\mathbf{r})/C_{0}^{(n)}(\mathbf{r})$ is bounded by its values for uniform distributions with $\theta_{k}=\theta_{\text{min}}$ and $\theta_{k}=\theta_{\text{max}}$ for all $k$, too. The derivative of $\bar{\theta}^{(n)}$, defined by Eq. (24), over the ratio $R$ has constant sign, independent of $R$:
$\frac{\partial\bar{\theta}^{(n)}(\mathbf{r})}{\partial R(\mathbf{r})}=\frac{A_{00}}{nA_{01}}\frac{(A_{10}/A_{00})^{n-1}A_{11}/A_{01}-(A_{10}/A_{00})^{n}}{[R(\mathbf{r})-(A_{10}/A_{00})^{n-1}A_{11}/A_{01}]^{2}}.$ (47)
Therefore, the values $\bar{\theta}^{(n)}(\mathbf{r})$ are also bounded:
$\Theta(\theta_{\text{min}})\leq\bar{\theta}^{(n)}(\mathbf{r})\leq\Theta(\theta_{\text{max}}).$ (48)
Finally, taking into account the definition (26) and monotonicity of $\Theta^{-1}$, we have
$\theta_{\text{min}}\leq\bar{\theta}^{(n)}_{\text{nonlin}}(\mathbf{r})\leq\theta_{\text{max}}.$ (49)
3. If the emitters are evenly distributed (i.e., for each point of interest $\mathbf{r}$ there is at least one emitter with $U(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}\geq C>0$ for some constant $C$), the parameter estimation by $\bar{\theta}^{(n)}_{\text{nonlin}}(\mathbf{r})$ is local:
$\frac{\partial\bar{\theta}^{(n)}_{\text{nonlin}}(\mathbf{r})}{\partial\theta_{k}}\rightarrow 0\text{ for }|\mathbf{r}-\mathbf{r}_{k}|\rightarrow\infty.$ (50)
The statement is valid if the PSF of the imaging system is local:
$U(|\mathbf{r}-\mathbf{r}_{k}|)\rightarrow 0\text{ for }|\mathbf{r}-\mathbf{r}_{k}|\rightarrow\infty.$ (51)
Then, the relative contribution of a distant emitter to the cumulant value becomes negligible for $|\mathbf{r}-\mathbf{r}_{k}|\rightarrow\infty$:
$\left|\frac{1}{C_{j}^{(n)}(\mathbf{r})}\frac{\partial C_{j}^{(n)}(\mathbf{r})}{\partial\theta_{k}}\right|\propto\frac{U^{n}(|\mathbf{r}-\mathbf{r}_{k}|)\varepsilon_{k}^{n}}{C^{n}}\rightarrow 0.$ (52)
Therefore, it does not influence the value of the estimator $\bar{\theta}^{(n)}_{\text{nonlin}}(\mathbf{r})$.
Appendix D Expressions for regularized parameter estimation
@@
$-\log p\left[\theta(\mathbf{r}_{ij})|\bar{\theta}(\mathbf{r}_{ij})\right]\propto\sum_{i,j}\frac{[\bar{\theta}(\mathbf{r}_{ij})-\theta(\mathbf{r}_{ij})]^{2}}{2\Delta\bar{\theta}^{2}(\mathbf{r}_{ij})}\\ +\sum_{i,j}\frac{[2\theta(\mathbf{r}_{ij})-\theta(\mathbf{r}_{i-1,j})-\theta(\mathbf{r}_{i+1,j})]^{2}}{2D^{2}}\\ +\sum_{i,j}\frac{[2\theta(\mathbf{r}_{ij})-\theta(\mathbf{r}_{i,j-1})-\theta(\mathbf{r}_{i,j+1})]^{2}}{2D^{2}}.$ (53)
@@
The sums representing the second and the third terms do not include boundary values of $i$ and $j$ respectively. The parameter $D$ scales as $D=const\times\langle\Delta\bar{\theta}(\mathbf{r}_{ij})\rangle w$, where $w$ is the PSF width expressed in the grid steps and representing the expected minimal size of the reconstructed distribution features, which are not artefacts yet. The proportionality coefficient can be calculated from detailed theoretical investigation of the noise models, but, in practice, it can be just chosen empirically.
@@
Unconstrained minimization of the quantity, defined by the right-hand side of Eq. (53), is trivial and includes just linear inversion according to the least-squares method. One of the approaches to obtaining physically reasonable constrained estimates $\theta_{\text{min}}\leq\theta(\mathbf{r}_{ij})\leq\theta_{\text{max}}$ is to perform constrained quadratic optimization, which, however, becomes much more computationally hard for a large number of grid points $\mathbf{r}_{ij}$. A more efficient way to get reasonable results is to constrain the initial estimate $\bar{\theta}(\textbf{r}_{ij})$ to the range $[\theta_{\text{min}},\theta_{\text{max}}]$ before applying the minimization of the expression (53). If all the values satisfy the constraint $\bar{\theta}(\textbf{r}_{ij})\in[\theta_{\text{min}},\theta_{\text{max}}]$, the minimizing set $\{\theta(\textbf{r}_{ij}\}$ will also belong to the physical range with certainty.

### v2

@@ V Conclusions
We believe that the proposed approach and its proof-of-concept experiment can pave the way for further research, including the following possible directions:
@@
  • extension of the approach to multimodal sensing, when several characteristics of the sample are reconstructed simultaneously;
@@ Acknowledgements.
This study was accomplished through the financial support of Research Council of Finland (Flagship Programme PREIN, decision 346518; research project, decision 368653); QuantERA project EXTRASENS, in Finland funded through Research Council of Finland, decision 361115), Horizon Europe MSCA FLORIN Project 101086142.
Appendix A Locally optimal signal compression
Here, we show that the information about the environment parameter $\theta$, contained in the normalized signal density profile $g(x)\equiv\tilde{f}(\mathbf{r}_{0},t,x)$, defined by Eq. (14), can be efficiently compressed into a ratio $Z$ of two scalar characteristics $Y_{i}$ (Eq. (15)) containing the same information. We assume that the signal profile is measured at a discrete set of bins $\{x\}$ and contains shot noise $\delta(x)$ satisfying Poisson statistics (i.e., we analyze shot noise for a fixed blinking realization since the blinking influences the in-frame signal profile as a whole):
$g(x)\mapsto g_{\text{noisy}}(x)=g(x)+\delta(x),$ (39)
where $\operatorname{E}[\delta(x)]=0$, $\operatorname{E}[\delta(x)\delta(x^{\prime})]\approx g(x)\delta_{xx^{\prime}}$. The high signal-to-noise ratio (SNR) regime is considered, i.e., $|\delta(x)|\ll|g(x)|$.
The Fisher information [76] of the normalized density profile $g(x)$ relatively to the parameter $\theta$ can be calculated as
$F_{\text{full}}(\theta)=\sum_{x}\frac{\left[g^{\prime}(x)\right]^{2}}{g(x)}-\frac{1}{\sum_{x}g(x)}\Bigl{(}\sum_{x}g^{\prime}(x)\Bigr{)}^{2},$ (40)
where the second term describes “information loss” caused by the normalization [77] and $g^{\prime}(x)=dg(x)/d\theta$.
Substitution of $g_{\text{noisy}}(x)$ into Eq. (15) and replacement of the integrals by sums yields
$Z\mapsto Z_{\text{noisy}}\\ \approx Z\left(1+\frac{\sum_{x}\delta(x)Q_{1}(x)}{\sum_{x}g(x)Q_{1}(x)}-\frac{\sum_{x}\delta(x)Q_{0}(x)}{\sum_{x}g(x)Q_{0}(x)}\right).$ (41)
In the considered high-SNR regime, $Z_{\text{noisy}}$ represents a normally distributed random variable with the expectation value $\operatorname{E}(Z_{\text{noisy}})=Z$ and the variance
$\operatorname{Var}(Z_{\text{noisy}})=\frac{1}{Y_{0}^{2}}\sum_{x}g(x)\bigl{[}Q_{1}(x)-Q_{0}(x)Z\bigr{]}^{2}.$ (42)
Therefore, Fisher information for the signal $Z_{\text{noisy}}$ relatively to the sensed parameter $\theta$ equals (see, e.g., Ref. [77])
$F_{Z}(\theta)=\frac{1}{\operatorname{Var}(Z_{\text{noisy}})}\left(\frac{dZ}{d\theta}\right)^{2}.$ (43)
According to general properties of Fisher information [78, 79], the compressed signal $Z_{\text{noisy}}$ cannot contain more information than the initial normalized profile $g_{\text{noisy}}(x)$ for any choice of the weights $Q_{0}(x)$ and $Q_{1}(x)$:
$F_{Z}(\theta)\leq F_{\text{full}}(\theta).$ (44)
On the other hand, by a specific choice of those weights, one can achieve equality of the two quantities. One can define
$Q_{0}(x)=\frac{q_{0}(x)}{\sum_{x^{\prime}}q_{0}(x^{\prime})g(x^{\prime})},\;Q_{1}(x)=\frac{q_{1}(x)}{\sum_{x^{\prime}}q_{1}(x^{\prime})g^{\prime}(x^{\prime})},$ (45)
@@ Appendix A Locally optimal signal compression
$q_{0}(x)=g(x)-g^{\prime}(x)\frac{\sum_{x^{\prime}}g(x^{\prime})g^{\prime}(x^{\prime})}{\sum_{x^{\prime}}\left[g^{\prime}(x^{\prime})\right]^{2}}$ (46)
@@
$q_{1}(x)=\frac{g^{\prime}(x)}{g(x)}-\frac{\sum_{x^{\prime}}g^{\prime}(x^{\prime})}{\sum_{x^{\prime}}g(x^{\prime})}.$ (47)
That specific choice of the weights ensures that
$Y_{0}=1,\quad Y_{1}=0,\quad\frac{dY_{0}}{d\theta}=0,\quad\frac{dY_{1}}{d\theta}=1$ (48)
and
$Z=0,\quad\frac{dZ}{d\theta}=1$ (49)
at the considered value of the parameter $\theta$ (the one used for construction of the weights). Substituting Eqs. (48) and (49) into Eqs. (42) and (43), we get the results $\operatorname{Var}(Z_{\text{noisy}})=1/F_{\text{full}}(\theta)$ and
$F_{Z}(\theta)=F_{\text{full}}(\theta).$ (50)
The choice of the weights is locally optimal. Generally, for values of $\theta$, other than the specific one used for the weights construction, the equality (50) may be violated. It is worth noting that in the high-SNR limit and in the vicinity of the considered parameter value $\theta$ the estimator
$\bar{\theta}=\zeta^{-1}(Z)\approx\theta+Z$ (51)
with the weights described by Eq. (45) coincidence with the locally unbiased estimator, saturating Cramer-Rao bound [80].
@@
Appendix B General properties of cumulant-based estimator for noiseless data
@@
To be physically reasonable, the estimator $\bar{\theta}^{(n)}(\mathbf{r})$, defined by Eq. (26) should satisfy several general requirements when applied to noiseless data. The requirements are provided by Eqs. (27)–(29). Here, we proof those statements.
1. For a spatially uniform distribution $\theta(\mathbf{r})=\theta_{0}$, one has $Y_{j}(\theta_{k})=Y_{j}(\theta_{0})$ for all emitter indices $k$. Therefore, according to Eq. (24),
$\frac{C_{1}^{(n)}(\mathbf{r})}{C_{0}^{(n)}(\mathbf{r})}=\frac{Y_{1}^{n}(\theta_{0})}{Y_{0}^{n}(\theta_{0})}=\zeta^{n}(\theta_{0})\text{ for all {\hbox{\mathbf{r}}}},$ (52)
and
$Z_{n}(\mathbf{r})=\zeta(\theta_{0})\quad\Rightarrow\quad\bar{\theta}^{(n)}(\mathbf{r})=\theta_{0}.$ (53)
2. If the parameter values $\theta_{k}$ in the vicinity of the emitters are bounded as $\theta_{k}\in[\theta_{\text{min}},\theta_{\text{max}}]$, the monotonicity assumption implies that the quantities $\zeta(\theta_{k})$ are also bounded as $\zeta(\theta_{k})\in[\zeta_{\text{min}},\zeta_{\text{max}}]$, where $\zeta_{\text{min}}=\min(\zeta(\theta_{\text{min}}),\zeta(\theta_{\text{max}}))$ and $\zeta_{\text{max}}=\max(\zeta(\theta_{\text{min}}),\zeta(\theta_{\text{max}}))$. Under the assumption that $\zeta_{\text{min}}\geq 0$ and $U(\mathbf{r}-\mathbf{r}_{k})\geq 0$, for each emitter, the following relation holds:
$Y_{1}^{n}(\theta_{k})\in Y_{0}^{n}(\theta_{k})\times[\zeta_{\text{min}}^{n},\zeta_{\text{max}}^{n}].$ (54)
Therefore, for any $\mathbf{r}$,
$C_{1}^{(n)}(\theta_{k})\in C_{0}^{(n)}(\theta_{k})\times[\zeta_{\text{min}}^{n},\zeta_{\text{max}}^{n}].$ (55)
Finally, $Z_{n}(\mathbf{r})\in[\zeta_{\text{min}},\zeta_{\text{max}}]$ and, therefore, $\bar{\theta}^{(n)}(\mathbf{r})\in[\theta_{\text{min}},\theta_{\text{max}}]$.
3. If the variations $\delta\theta_{k}=\theta_{k}-\theta_{0}$ are small, one can use the linear approximation and get
$Y_{j}^{n}(\theta_{k})\approx Y_{j}^{n}(\theta_{0})+nY_{j}^{n-1}(\theta_{0})Y_{j}^{\prime}(\theta_{0})\delta\theta_{k},$ (56)
[…]
$C_{j}^{(n)}(\mathbf{r})\approx Y_{j}^{n}(\theta_{0})\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\\ +nY_{j}^{n-1}(\theta_{0})Y_{j}^{\prime}(\theta_{0})\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\delta\theta_{k}$ (57)
and
$Z_{n}(\mathbf{r})\approx\zeta(\theta_{0})\Biggl{[}1+\left(\frac{Y_{1}^{\prime}(\theta_{0})}{Y_{1}(\theta_{0})}-\frac{Y_{0}^{\prime}(\theta_{0})}{Y_{0}(\theta_{0})}\right)\\ \times\frac{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\delta\theta_{k}}{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]}\Biggr{]}$ (58)
Since
$\zeta(\theta_{0}+\delta\theta)\approx\zeta(\theta_{0})\left[1+\left(\frac{Y_{1}^{\prime}(\theta_{0})}{Y_{1}(\theta_{0})}-\frac{Y_{0}^{\prime}(\theta_{0})}{Y_{0}(\theta_{0})}\right)\delta\theta\right],$ (59)
one can easily find the solution of the equation $\zeta(\theta)=Z_{n}(\mathbf{r})$ in the linear approximation and conclude that
$\bar{\theta}^{(n)}(\mathbf{r})=\theta_{0}+\frac{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\delta\theta_{k}}{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]}\\ =\frac{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]\theta_{k}}{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]}.$ (60)
Appendix C Expressions for regularized parameter estimation
@@
$-\log p\left[\theta(\mathbf{r}_{ij})|\bar{\theta}^{(n)}(\mathbf{r}_{ij})\right]\propto\sum_{i,j}\frac{[\bar{\theta}^{(n)}(\mathbf{r}_{ij})-\theta(\mathbf{r}_{ij})]^{2}}{2\left[\Delta\bar{\theta}^{(n)}(\mathbf{r}_{ij})\right]^{2}}\\ +\sum_{i,j}\frac{[2\theta(\mathbf{r}_{ij})-\theta(\mathbf{r}_{i-1,j})-\theta(\mathbf{r}_{i+1,j})]^{2}}{2D^{2}}\\ +\sum_{i,j}\frac{[2\theta(\mathbf{r}_{ij})-\theta(\mathbf{r}_{i,j-1})-\theta(\mathbf{r}_{i,j+1})]^{2}}{2D^{2}}.$ (61)
@@
The sums representing the second and the third terms do not include boundary values of $i$ and $j$ respectively. The parameter $D$ scales as $D=\text{const}\times\langle\Delta\bar{\theta}^{(n)}(\mathbf{r}_{ij})\rangle w$, where $w$ is the PSF width expressed in the grid steps and representing the expected minimal size of the reconstructed distribution features, which are not artifacts yet. The proportionality coefficient can be calculated from detailed theoretical investigation of the noise models, but, in practice, it can be just chosen empirically.
@@
Unconstrained minimization of the quantity, defined by the right-hand side of Eq. (61), is trivial and consists in linear inversion according to the least-squares method. One of the approaches to obtaining physically reasonable constrained estimates $\theta_{\text{min}}\leq\theta(\mathbf{r}_{ij})\leq\theta_{\text{max}}$ is to perform constrained quadratic optimization, which, however, becomes much more computationally hard for a large number of grid points $\mathbf{r}_{ij}$. A more efficient way to get reasonable results is to constrain the initial estimate $\bar{\theta}^{(n)}(\textbf{r}_{ij})$ to the range $[\theta_{\text{min}},\theta_{\text{max}}]$ before applying the minimization of the expression (61). If all the values satisfy the constraint $\bar{\theta}^{(n)}(\textbf{r}_{ij})\in[\theta_{\text{min}},\theta_{\text{max}}]$, the minimizing set $\{\theta(\textbf{r}_{ij}\}$ will also belong to the physical range with certainty.
Appendix D SOFI with cross-cumulants
The expression (10) for a the zero time lag cross-cumulant
$C_{t}^{(n)}[F(\mathbf{r}^{(1)},t),\ldots,F(\mathbf{r}^{(n)},t)]\\ \propto\sum_{k}U^{n}(\bar{\mathbf{r}}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)],$ (62)
coincides with Eq. (4) for an autocumulant $C_{t}^{(n)}[F(\bar{\mathbf{r}},t)]$ up to a constant multiplier. Therefore, to improve SNR, one can replace the autocumulant signal $C_{t}^{(n)}(\mathbf{r})$ used in SOFI by a weighted combination of cross-cumulants:
$C_{t}^{(n)}(\mathbf{r})\mapsto\tilde{C}_{t}^{(n)}(\mathbf{r})=\sum_{\begin{array}[]{c}\scriptstyle\mathbf{r}^{(1)},\ldots,\mathbf{r}^{(n)}:\\ \scriptstyle\mathbf{r}^{(1)}+\cdots+\mathbf{r}^{(n)}=n\mathbf{r}\end{array}}w_{\mathbf{r}^{(1)},\ldots,\mathbf{r}^{(n)}}\\ \times C_{t}^{(n)}[F(\mathbf{r}^{(1)},t),\ldots,F(\mathbf{r}^{(n)},t)].$ (63)
During processing of the simulated data, we choose the weights as
$w_{\mathbf{r}^{(1)},\ldots,\mathbf{r}^{(n)}}\propto\exp\left(-\frac{\sum_{j}\left|\mathbf{r}^{(j)}\right|^{2}}{\sigma^{2}}\right)$ (64)
with $\sigma=3\delta$, where $\delta$ is the detection step (detector pixel size, recalculated to the object plane of the microscope). As a compromise between noise reduction and computational complexity, we additionally restrict the sum in Eq. (63) by the condition $\sum_{j}\left|\mathbf{r}^{(j)}\right|^{2}\leq 5\sigma^{2}$.
The approach can also be extended to non-zero time lag cross-cumulants, but we do not consider such modification since the goal of simulations is illustration of the proposed sensing approach rather than achievement of the ultimate accuracy of SOFI results.
The source code of the described method implementation is available in Supplemental Material at [URL will be inserted by publisher].
Appendix E Influence of PSF shape on resolution enhancement
According to Eq. (4), the resolution enhancement in SOFI is provided by raising PSF $U(\Delta\mathbf{r})$ to the power $n$ for the $n$-th order cumulant image. If the PSF can be approximated by a Gaussian profile,
$U(\Delta\mathbf{r})\propto\exp\left(-\frac{\Delta\mathbf{r}^{2}}{2w^{2}}\right),$ (65)
it preserves the shape after raising to the power $n$ and acquires $\sqrt{n}$ times smaller width (Fig. 13(a)):
$U^{n}(\Delta\mathbf{r})\propto\exp\left(-n\frac{\Delta\mathbf{r}^{2}}{2w^{2}}\right)=\exp\left(-\frac{\Delta\mathbf{r}^{2}}{2w_{n}^{2}}\right),$ (66)
where $w_{n}=w/\sqrt{n}$. Similar scaling is characteristic for the Airy disk profile (so called “sombrero” function [81]), obtained for the blurring caused by diffraction on a lens with finite-radius aperture (Fig. 13(b)). Such behavior of typical PSF shapes allows one to state that $n$-th order SOFI enhances resolution by the factor $\sqrt{n}$.
However, the behavior can be significantly different for other PSF shapes. In the limiting case of PSF being the rectangular function
$U(\Delta\mathbf{r})\propto\operatorname{rect}(|\Delta\mathbf{r}|/w)=\left[\begin{array}[]{cc}1&|\Delta\mathbf{r}|<w,\\ 0&\text{otherwise},\end{array}\right.$ (67)
raising it to any power $n>0$ will not influence the blurring width (Fig. 13(c)):
$U^{n}(\Delta\mathbf{r})\propto U(\Delta\mathbf{r}).$ (68)
In the intermediate case of convolved Gaussian and rectangular profiles,
$U(\Delta\mathbf{r})\propto\exp\left(-\frac{\Delta\mathbf{r}^{2}}{2w_{\text{G}}^{2}}\right)*\operatorname{rect}(|\Delta\mathbf{r}|/w_{\text{R}}),$ (69)
the resolution enhancement scales with the growth of $n$ much slower than $\sqrt{n}$ (almost saturates, Fig. 13(d, e)).
Figure 13: Influence of PSF shape on resolution enhancement. (a-d) Profiles of PSF (with argument normalized by full width at half maximum — FWHM) raised to the power $n=1$ (blue solid lines), 2 (yellow dashed lines), 4 (greed dotted lines), and 8 (red dot-dashed lines) for Gaussian (a), “sombrero” (b), rectangular (c) profiles, and the convolution of Gaussian and rectangular profiles (d) with $w_{\text{G}}:w_{\text{R}}=1:10$. (e) Normalized FWHM of PSF raised to the power $n$ for for Gaussian (blue solid lines), “sombrero” (yellow dashed lines), rectangular (greed dotted lines) profiles, and the convolution of Gaussian and rectangular profiles (red dot-dashed lines).